A uniform field of 2.0 N/C exists in space in x-direction. (a) Taking the potential at the origin to be zero, write an expression for the potential at a general point (x, y, z). (b) At which points, the potential is 25 V? (c) If the potential at the origin is taken to be 100 V, what will be the expression for the potential at a general point? (d) What will be the potential at the origin if the potential at infinity is taken to be zero? Is it practical to choose the potential at infinity to be zero?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. E = 2 N/C in x-direction (a) Potential aat the origin is O. dV = – E base x dx – E base y dy – E base z dz ⇒ V – 0 = – 2x ⇒ V = – 2x (b) (25 – 0) = – 2x ⇒ x = – 12.5 m (c) If potential at origin is 100 v, v – 100 = – 2x ⇒ V = – 2x + 100 = 100 – 2x (d) Potential at ∞ IS 0, V – V’ = – 2x ⇒ V’ = V + 2x = 0 + 2∞ ⇒ V’ = ∞ Potential at origin is ∞. No, it is not practical to take potential at ∞ to be zero.